Desa invested a certain amount of money at 2% interest and $12000 more than that amount at 7%. Her total yearly interest was $545. How much did she invest at each rate?

Answer by stanbon(75887)

Desa invested a certain amount of money at 2% interest and $12000 more than that amount at 7%. Her total yearly interest was $545. How much did she invest at each rate?
-------------------
Equation::
interest + interest = interest
-----
0.02x + 0.07(12000-x) = 545
------
2x + 7*12000 - 7x = 54500
------
-5x = -29500
x = $5900 (amt. invested at 2%)
12000-x = 12000-5900 = $6100 (amt. invested at 7%)
--------------
Cheers,
Stan H.
--------------